# LightGaussian: Unbounded 3D Gaussian Compression with 15x Reduction and 200+ FPS

### v5

@@ 3.4 Gaussian Attributes Vector Quantization
Vector quantization clusters the voxels into compact codebooks. While high compression rates are attainable, the task of quantizing Gaussian attributes in a point-based representation, inherently non-Euclidean, presents notable challenges. Our empirical investigations reveal that quantization applied to all elements, particularly attributes such as position, rotation, and scale, tends to incur significant accuracy losses when represented discretely, leading to a marked reduction in precision.
Vector Quantization.
We propose applying Vector Quantization (VQ) to the Spherical Harmonics (SH) coefficients, based on the assumption that a subgroup of 3d Gaussians typically exhibits a similar appearance. Fundamentally, VQ segments the Gaussians $\mathcal{G}=\{\mathbf{g}_{1},\mathbf{g}_{2},\ldots,\mathbf{g}_{N}\}$ (here we apply it on SH) to the K codes in the codebook $\mathcal{C}=\{\mathbf{c}_{1},\mathbf{c}_{2},\ldots,\mathbf{c}_{K}\}$, where each $\mathbf{g}_{j},\mathbf{c}_{k}\in\mathbb{R}^{d}$ and K $\ll$ N. $d$ means SH dimension. We aim to strike a balance between rendering quality loss and compression rate by leveraging the pre-computed significance score from Eq. 3. Based on this, we apply Vector Quantization (VQ) selectively on the least significant elements in the Spherical Harmonics (SHs). Specifically, we initialize $\mathcal{C}$ via K-means, iteratively sample a batch of $\mathcal{G}$, associates them to the closest codes by euclidean distance, and update each $\mathbf{c}_{k}$ via moving average rule: $\mathbf{c}_{k}=\lambda_{d}\cdot\mathbf{c}_{k}+(1-\lambda_{d})\cdot 1/\mathbf{T}_{k}\cdot\sum_{\mathbf{g}_{j}\in\mathcal{R}(\mathbf{c}_{k})}\mathbf{GS}_{j}\cdot\mathbf{g}_{j}$, where $\mathbf{T}_{k}=\sum_{\mathbf{g}_{j}\in\mathcal{R}(\mathbf{c}_{k})}{\mathbf{GS}_{j}}$ is the significance score (Eq. 3) which is assigned to the code vector $\mathbf{c}_{k}$, $\mathcal{R}(\mathbf{c}_{k})$ is the set of Gaussians associated to the $k$-th code. $\lambda_{d}=0.8$ represents the decay value, which is utilized to update the code vector using a moving average. We fine-tune the codebook for 5,000 iterations, while fixing the gaussian-to-codebook mapping. We disable additional clone/split operations and leverage photometric loss on the training views.
To maintain the integrity of crucial attributes such as these Spherical Harmonics with higher global significance scores, and Gaussian’s position, shape, rotation, and opacity attributes, we opt to bypass the VQ process and instead store these features directly in the float16 format.
Table 1: Quantitative Comparisons in Real-world Large-scale Scenes. We compare LightGaussian with the original 3D-GS [32], efficient voxel-based NeRFs [67, 46], voxel NeRF with vector quantization [40], and compact MLP-based NeRF [5]. Voxel-based methods all exhibit a lack of sufficient capacity for representing large-scale scenes and are not able to run at real-time speed. Mip-NeRF 360 produces the best visual quality, but it requires more than 16s to render a single image. Our method achieves a good balance among FPS, model size, and rendering quality. We re-train 3D-GS using their original code and configurations on our platform, to perform a fair comparison, noted as 3D-GS.
@@
3D-GS [32] 119 724MB 27.53 0.810 0.227 106 380MB 23.11 0.822 0.219
Ours 209 42MB 27.28 0.805 0.243 209 22MB 23.11 0.817 0.231
Figure 4: Visual Comparisons. We compare LightGaussian with the vanilla 3D-GS [32], presenting a residual map between prediction and ground-truth scaled from 0 to 127 to highlight the differences. We observe that LightGaussian could preserve most of the specular reflection (yellow boxes) after converting to a compact format. Additionally, a slight lightness change is noted after the conversion, as shown in the bottom white box. Note that the residual maps plotted illustrate the differences between the rendered images and the GT RGB values. Here, darker regions indicate closer alignment. For a comprehensive dynamic viewpoint comparison, please refer to our supplementary video material.
@@ 4.1 Experimental Settings
We conduct comparisons using the scene-scale view synthesis dataset provided by Mip-NeRF360 [6], which comprises nine real-world large-scale scenes, including five unbounded outdoor and four indoor settings with intricate backgrounds. Furthermore, we utilize the Tanks and Temples dataset [34], a comprehensive unbounded dataset, for additional comparisons, selecting the same scenes as those used in 3D-GS [32]. Performance metrics on synthetic object-level datasets will be detailed in the supplementary materials due to constraints on space. We report metrics including the peak signal-to-noise ratio (PSNR), structural similarity (SSIM), and perceptual similarity as measured by LPIPS [70].
@@
We conduct comparisons with methodologies capable of modeling large-scale scenes. This includes Plenoxel [67], Mip-NeRF360 [5], and 3D-GS [32]. Additionally, we compare against efficient methods like Instant-ngp [46], which leverages a hash grid for storage efficiency, and VQ-DVGO [40]. The latter builds upon the DVGO framework [57], incorporating enhancements such as voxel pruning and vector quantization for optimized representation.
Table 2: Ablation studies on the Gaussian Pruning & Recovery, SH Compactness, and the Vector Quantization. Scene: Room. Zero-shot Gaussian pruning leads to a degradation in rendering quality (#2), but Co-adaptation can recover most of the scene details (#3). Directly eliminating high-order SH negatively affects the quality (#4), while distillation with pseudo-view helps to mitigate the gap (#5, #6). Codebook quantization further significantly reduces the required model size and bandwidth (#7).
@@
[1] Baseline (3D-GS [32]) 192.05 353MB 31.68 0.926 0.200
[2]    + Gaussian Pruning. 312.30 116MB 30.32 0.911 0.222
[3]     + Co-adaptation 303.99 116MB 31.85 0.925 0.206
[4]    + SH Compactness. 318.97 77MB 30.54 0.914 0.217
[5]     + Distillation 304.20 77MB 31.47 0.922 0.211
[6]     + Pseudo-views 300.60 77MB 31.59 0.923 0.211
[7]    + Codebook Quant. 300.60 20MB 31.16 0.915 0.218
[8]    + VQ finetune 300.60 20MB 31.67 0.923 0.212
[9] LightGaussian (Ours) 300.60 20MB 31.67 0.923 0.212
@@
Baseline 192.05 353MB 31.68 0.926 0.200
Hit Count Only 300.57 116MB 28.28 0.895 0.238
   $\times$ Opacity. 310.29 116MB 30.10 0.910 0.222
   $\times$ Opacity $\times$ $\gamma$(Volume). 312.30 116MB 30.32 0.911 0.222
    + Co-adaptation 303.99 116MB 31.85 0.925 0.206
Table 4: Ablation study Gaussian Attribute Vector Quantization (VQ). By quantizing all attributes to FP16, a smaller model is achieved. VQ is applied to all attributes leads to inferior modeling accuracy, but this can be mitigated by using Global Significance (GS) on the least crucial Gaussians. Other attributes (e.g., scale) are also sensitive to VQ, hence we only apply VQ on the SH. Combining VQ and significance score as indicators, LightGaussian demonstrates a good balance between size and quality.
[…]
Baseline 77.09MB 31.59 0.923 0.211
+FP16 33.56MB 31.47 0.921 0.212
   + VQ All att. 18.74MB 23.02 0.748 0.376
   + VQ All att. $\times$ GS 18.74MB 26.33 0.831 0.325
   + VQ SH. 19.89MB 30.85 0.910 0.222
   + VQ SH $\times$ GS 19.89MB 31.16 0.915 0.218
LightGaussian (w/ VQ Finetune) 19.89MB 31.67 0.923 0.212
@@
Our framework is implemented using Pytorch and incorporates the differential Gaussian rasterization technique from 3D-GS [32]. We utilize our own trained models, ensuring they align with the performance benchmarks established in the 3D-GS study [32]. All performance are evaluated on a A6000 GPU, the same as 3D-GS. In the Global Significance Calculation phase, we assign a power value of 0.1 in Eq. 4 and proceed to fine-tune the model for 5,000 steps during the Gaussian Co-adaptation process. For SH Distillation, we downscale the 3-degree SHs to 2-degree, thereby reducing 21 elements for each Gaussian. This process is further optimized by setting $\sigma$ to 0.1 in the pseudo view synthesis stage. In the Gaussian Vector Quantization step, the codebook size is configured to 8192, selecting Spherical Harmonics (SHs) with the least $60\%$ significance score for the vector quantization (VQ ratio), to balance the trade-off between compression efficiency and fidelity.
Figure 5: Visual Comparisons for Ablation Study. We visualize the rendered RGB images and the residual map between the ground-truth image, aligned with the experiment ID as shown in Tab. 2. The final model (Exp #9) demonstrates close results to 3D-GS (Exp #1), while the Gaussian Co-adaptation, along with SH distillation, almost completely mitigates the information loss. To highlight the difference, we visualize the SSIM between the rendered images and the GT, where a whiter output in the residual signifies closer alignment with the GT, denoting better quality, and the area with greater color intensity indicates lower quality.
@@ 4.2 Experimental Results
We evaluate the performance of various methods for novel view synthesis by summarizing their quantitative metrics in Tab. 7. This includes highly efficient voxel-based NeRFs such as Plenoxel[67] and Instant-NGP [46], the compact MLP-based Mip-NeRF360 [5], NeRF enhanced with vector quantization [40], and 3D Gaussian Splatting [32]. Specifically, within the Mip-NeRF360 dataset, NeRF-based methods leverage MLPs for a compact representation, achieving competitive accuracy. However, their slow inference speed (0.06FPS) limits practical applications. Voxel-based NeRFs improve rendering efficiency but still fall short of real-time performance. For instance, Plenoxel demands 2.1GB for storing a large-scale scene. VQ-DVGO addresses storage concerns yet faces generalization issues in complex large-scale scenes due to its reliance on pruning and vector quantization alone. Conversely, 3D-GS offers a commendable balance between rendering quality and speed but requires substantial storage for a single scene. Our method, LightGaussian, surpasses existing techniques with a rendering speed exceeding 200 FPS, courtesy of its efficient rasterization process that prunes insignificant Gaussians. Collectively, the proposed techniques significantly reduces the redundancy in 3D Gaussian model, cutting down the size from 724MB to 42MB on the Mip-NeRF360 datasets—a reduction factor of 15$\times$. LightGaussian also outperforms 3D-GS in rendering speed on the Tank & Temple datasets by nearly 2$\times$ while shrinking the storage requirement from 380MB to 22MB.
@@
We conducted a comparative analysis of the rendering outcomes between 3D-GS and our LightGaussian model, focusing particularly on the intricate details and background regions, as demonstrated in Fig. 4. It is evident that both 3D-GS and LightGaussian maintain comparable visual quality, even in scenes with thin structures (refer to the first and third rows).
@@ 4.3 Ablation Studies
We conducted ablation studies on our method’s components to individually assess the impact of the proposed modules. Our findings reveal that Gaussian Pruning and Recovery is effective in eliminating redundant Gaussians, ensuring that only significant ones contribute to the scene representation. SH Distillation proves to be effective in reducing the degree of Spherical Harmonics, which simplifies the lighting model without compromising too much quality. Additionally, Vector Quantization demonstrates its efficiency in compressing the feature space, resulting in a more streamlined and efficient representation. When combined, these modules significantly enhance the overall efficacy and performance of our framework.
Overall Analysis.
@@
The Best Significance Criteria?
To discover the most favorable criteria to measure the global significance of each Gaussian, we evaluated key characteristics within each Gaussian: the Gaussian/rays interaction count, Gaussian opacity, and the volume related to its function. Tab. 3 illustrates that integrating these three factors—by weighting the hit count with opacity and Gaussian volume—yields the highest rendering quality post zero-shot pruning. A quick Gaussian Co-adaptation (the last row) which optimizes the remaining significant Gaussians, can effectively recover the rendering accuracy to its pre-pruning level. The visual comparison of rasterized images, before and after pruning, alongside the depiction of pruned Gaussians, is presented in Fig. 6.
@@
The direct removal of high-degree components in Spherical Harmonics (SHs) results in a substantial performance decline compared to the full model (Exp #3 $\rightarrow$ 4), manifesting notably in the loss of specular reflections on surfaces with varying viewpoints. However, the integration of knowledge distillation (Exp #4 $\rightarrow$ 5) from the full model enables us to decrease its size while maintaining the essential viewing effects. Furthermore, the incorporation of pseudo-views during the training process (Exp #5 $\rightarrow$ 6) validate the efficacy of using the simulated views to enhance the learning of specular reflections. Applying Vector Quantization (VQ) on all attributes leads to inferior results (See Tab. 4), but this can be mitigated by using Gaussian Global Significance.
Table 5: Ablation under different SH ompactness: With a Gaussian pruning ratio of 66% and reduce SH to 2-degree, LightGaussian nearly preserves the rendering quality (SSIM decrease from 0.927 to 0.926). Reducing the SH to 1-degree lead to the SSIM decreases to 0.923.
@@
Figure 7: Visualization of the rendering performance vs. different prune and VQ ratios. The quantize ratio is the compression ratio at vector quantization.
@@
5 More Technical Details

### v6

@@ 3.4 Gaussian Attributes Vector Quantization
Vector quantization (VQ) clusters voxels into compact codebooks, enabling high compression rates. However, quantizing Gaussian attributes in a point-based, inherently non-Euclidean representation poses notable challenges. Our empirical findings indicate that applying quantization across all elements—especially for attributes like position, rotation, and scale—results in significant accuracy losses and a marked reduction in precision when represented discretely.
We propose applying VQ to the Spherical Harmonics (SH) coefficients, based on the assumption that a subgroup of 3d Gaussians typically exhibits a similar appearance. Fundamentally, VQ segments the Gaussians $\mathcal{G}=\{\mathbf{g}_{1},\mathbf{g}_{2},\ldots,\mathbf{g}_{N}\}$ (here we apply it on SH) to the K codes in the codebook $\mathcal{C}=\{\mathbf{c}_{1},\mathbf{c}_{2},\ldots,\mathbf{c}_{K}\}$, where each $\mathbf{g}_{j},\mathbf{c}_{k}\in\mathbb{R}^{d}$ and K $\ll$ N. $d$ means SH dimension. We aim to strike a balance between rendering quality loss and compression rate by leveraging the pre-computed significance score from Eq. 3. Based on this, we apply VQ selectively on the least significant elements in the Spherical Harmonics (SHs). Specifically, we initialize $\mathcal{C}$ via K-means, iteratively sample a batch of $\mathcal{G}$, associates them to the closest codes by euclidean distance, and update each $\mathbf{c}_{k}$ via moving average rule: $\mathbf{c}_{k}=\lambda_{d}\cdot\mathbf{c}_{k}+(1-\lambda_{d})\cdot 1/\mathbf{T}_{k}\cdot\sum_{\mathbf{g}_{j}\in\mathcal{R}(\mathbf{c}_{k})}\mathbf{GS}_{j}\cdot\mathbf{g}_{j}$, where $\mathbf{T}_{k}=\sum_{\mathbf{g}_{j}\in\mathcal{R}(\mathbf{c}_{k})}{\mathbf{GS}_{j}}$ is the significance score (Eq. 3) which is assigned to the code vector $\mathbf{c}_{k}$, $\mathcal{R}(\mathbf{c}_{k})$ is the set of Gaussians associated to the $k$-th code. $\lambda_{d}=0.8$ represents the decay value, which is utilized to update the code vector using a moving average. We fine-tune the codebook for 5,000 iterations, while fixing the gaussian-to-codebook mapping. We disable additional clone/split operations and leverage photometric loss on the training views. To preserve essential attributes, including Spherical Harmonics with higher global significance scores, along with Gaussian position, shape, rotation, and opacity, we skip VQ for these elements and store them directly in float16 format.
@@
3D-GS [16] 144 782MB 27.40 0.813 0.217 106 433MB 23.66 0.845 0.178
Ours 237 45MB 27.13 0.806 0.237 357 25MB 23.44 0.832 0.202
Table 1: Quantitative Comparisons in Real-world Large-scale Scenes: Voxel-based methods [12, 14] exhibit insufficient capacity for representing large-scale scenes and are unable to achieve real-time performance. Mip-NeRF360 [7] produces the highest visual quality, but requires over 16 seconds to render a single image. Our method strikes a balance among FPS, model size, and rendering quality, achieving the best balance among efficient representations. For fair metric and visual comparison, we re-trained 3D-GS and Compressed 3D-GS [34] using their original code and configurations on our platform (NVIDIA A6000 GPU), with results marked by .
Figure 4: Visual Comparisons. We compare LightGaussian with the vanilla 3D-GS [16], displaying a residual map between predictions and ground truth scaled from 0 to 127 to emphasize differences. LightGaussian retains most specular reflections (yellow boxes) in its compact format, with a slight change in lightness visible in the bottom white box. The residual maps illustrate discrepancies between rendered images and ground-truth RGB values, where darker areas indicate closer alignment. For a full dynamic viewpoint comparison, please see our supplementary video.
@@ 4.1 Experimental Settings
We conduct comparisons using the scene-scale view synthesis dataset provided by Mip-NeRF360 [64], which comprises nine real-world large-scale scenes, including five unbounded outdoor and four indoor settings with complex backgrounds. In addition, we utilize the Tanks and Temples dataset [65], a comprehensive unbounded dataset, and select the same scenes as used in [16]. Performance metrics on synthetic object-level datasets will be detailed in the supplementary materials. We report metrics including the peak signal-to-noise ratio (PSNR), structural similarity (SSIM), and perceptual similarity as measured by LPIPS [66].
@@
We compare our approach with methods suited for large-scale scene modeling, including Plenoxel [12], Mip-NeRF360 [7], and 3D-GS [16]. Additionally, we evaluate against efficient techniques like Instant-NGP [14], which uses a hash grid for storage efficiency, and VQ-DVGO [49], which extends DVGO [9] with voxel pruning and VQ for optimized representation.
@@
[1] Baseline (3D-GS [16]) 156.21 365MB 31.34 0.917 0.221
[2]    + Gaussian Pruning 308.10 123MB 30.67 0.910 0.234
[3]     + Co-adaptation 304.64 123MB 31.64 0.918 0.228
[4]    + SH Compactness 311.58 81MB 30.32 0.904 0.238
[5]     + photometric loss 302.20 81MB 31.42 0.916 0.234
[6]     + Distillation + Pseudo-views 302.89 81MB 31.48 0.917 0.231
[7]    + Codebook Quant. 301.21 21MB 31.09 0.918 0.236
[8]    + VQ finetune 302.01 21MB 31.40 0.916 0.232
[9] LightGaussian (Ours) 302.01 21MB 31.40 0.916 0.232
Table 2: Ablation studies on the Gaussian Pruning & Recovery, SH Compactness, and the Vector Quantization. Scene: Room. Zero-shot Gaussian pruning leads degraded rendering quality (#2), but Co-adaptation can recover most of the scene details (#3). Directly eliminating high-order SH negatively affects the quality (#4), while distillation with pseudo-view helps to mitigate the gap (#5, #6). Codebook quantization further reduces the required model size and bandwidth (#7, #8).
[…]
Baseline 156.21 365MB 31.34 0.917 0.221
Hit Count Only 301.52 123MB 28.16 0.886 0.261
    + Co-adaptation 303.43 123MB 30.13 0.912 0.238
   $\times$ Opacity 310.29 123MB 30.27 0.909 0.239
    + Co-adaptation 304.84 123MB 31.60 0.916 0.231
   $\times$ Opacity $\times$ $\gamma$(Volume). 312.30 123MB 30.67 0.910 0.234
    + Co-adaptation 304.64 123MB 31.64 0.918 0.228
@@
Baseline 80.99MB 31.48 0.917 0.231
+FP16 36.51MB 31.35 0.914 0.232
   + VQ All att. 18.74MB 23.11 0.731 0.378
   + VQ All att. $\times$ GS 19.74MB 26.23 0.826 0.323
   + VQ SH. 21.10MB 30.68 0.907 0.244
   + VQ SH $\times$ GS 21.10MB 31.16 0.911 0.235
LightGaussian (w/ VQ Finetune) 21.10MB 31.40 0.916 0.232
Table 4: Ablation Study: Gaussian Attribute Vector Quantization (VQ). Quantizing all attributes to FP16 achieves a smaller model, but applying VQ to all attributes degrades modeling accuracy. Using Global Significance (GS) to target less crucial Gaussians mitigates this loss. Some attributes (e.g., scale) are sensitive to VQ, so we limit VQ to the SH coefficients. By combining VQ with significance scores, LightGaussian achieves an effective balance between model size and quality.
@@
Our framework is implemented in PyTorch and integrates the differentiable Gaussian rasterization technique from 3D-GS [16]. All performance evaluations are conducted on an A6000 GPU. In the Global Significance Calculation phase, we assign a power value of 0.1 in Eq. 4 and proceed to fine-tune the model for 5,000 steps during the Gaussian Co-adaptation process. For SH distillation, we downscale the 3-degree SHs to 2-degree, thereby reducing 21 elements for each Gaussian. This is further optimized by setting $\sigma$ to 0.1 in the pseudo view synthesis stage. In the Gaussian VQ step, the codebook size is configured to 8192, selecting SHs with the least $60\%$ significance score for the vector quantization (VQ ratio), to balance the trade-off between compression efficiency and fidelity.
@@ 4.2 Experimental Results
We assess the performance of various methods for novel view synthesis, with quantitative metrics summarized in Tab. 1. This includes efficient voxel-based NeRFs like Plenoxel [12] and Instant-NGP [14], the compact MLP-based Mip-NeRF360 [7], vector-quantized NeRF [49], and 3D Gaussian Splatting [16].
On the Mip-NeRF360 dataset, MLP-based NeRF methods achieve competitive accuracy with compact representations but suffer from slow inference speeds (0.06 FPS), limiting their practicality. Voxel-based NeRFs improve rendering efficiency but still fall short of real-time performance; for example, Plenoxel requires 2.1GB for a single large-scale scene. In contrast, 3D-GS offers a good balance between quality and speed but demands substantial storage per scene.
Our method, LightGaussian, exceeds existing techniques with rendering speeds over 200 FPS, enabled by efficient rasterization that prunes insignificant Gaussians. This approach reduces 3D Gaussian model redundancy, cutting storage from 782MB to 45MB on Mip-NeRF360—a 15$\times$ reduction. LightGaussian also outperforms 3D-GS on the Tank & Temple datasets, nearly doubling rendering speed and reducing storage from 380MB to 22MB.
@@
We conducted a comparative analysis of rendering results between 3D-GS and LightGaussian, focusing on intricate details and background regions, as shown in Fig. 4. Both 3D-GS and LightGaussian exhibit comparable visual quality, even in challenging scenes with thin structures, showing that LightGaussian effectively removes redundancy while preserving reconstruction fidelity.
Generalization to Other Point-based Representations
We further apply our Gaussian Pruning approach to Scaffold-GS [67], an advanced neural Gaussian-based 3D representation. In this method, we prune Gaussians that contribute minimally to scene reconstruction, reducing redundancy while preserving visual fidelity. Experiments on the MipNeRF360 dataset demonstrate that pruning 80% of neural Gaussians increases the rendering speed from 152 to 178 FPS, as shown in Tab. 5. These results underscore LightGaussian’s potential as an effective optimization tool for other Gaussian-based representations.
Size FPS PSNR SSIM LPIPS
Scaffold-GS 173.60 152 27.96 0.8240 0.2075
Scaffold-GS + LightGaussian 112.56 178 27.78 0.8187 0.2197
Table 5: LightGaussian removes redundant neural Gaussians in Scaffold-GS, reducing model size and improving rendering speed. All experiments were rerun on our platform for fair comparison, with results averaged on the MipNeRF-360 dataset.
@@ 4.3 Ablation Studies
We performed ablation studies on each component of our method to evaluate their individual impacts. Results show that Gaussian Pruning and Recovery effectively removes redundant Gaussians, retaining only those crucial for scene representation. SH Distillation reduces the Spherical Harmonics degree, simplifying the lighting model with minimal quality loss. Furthermore, Vector Quantization efficiently compresses the feature space, creating a more compact representation. Combined, these modules substantially improve the overall efficiency and performance of our framework.
Significance Criteria
To identify the optimal criteria for measuring each Gaussian’s global significance, we evaluated key characteristics: Gaussian-ray interaction count, Gaussian opacity, and functional volume. Tab. 3 illustrates that integrating these three factors — by weighting the hit count with opacity and Gaussian volume — yields the highest rendering quality post zero-shot pruning. A quick Gaussian Co-adaptation (the last row) which optimizes the remaining Gaussians, can effectively recover the rendering accuracy to its pre-pruning level. The visual comparison of rasterized images, before and after pruning, alongside the depiction of pruned Gaussians, is presented in Fig. 5.
@@
Directly removing high-degree components in Spherical Harmonics (SHs) causes significant performance degradation compared to the full model (Exp #3 $\rightarrow$ 4), particularly with the loss of specular reflections across varying viewpoints. However, introducing knowledge distillation from the full model (Exp #4 $\rightarrow$ 5) allows for size reduction while preserving essential viewing effects. Additionally, incorporating pseudo-views during training (Exp #5 $\rightarrow$ 6) demonstrates the effectiveness of simulated views in enhancing the learning of specular reflections. While applying VQ to all attributes yields poorer results (see Tab. 4), this impact is mitigated by leveraging Gaussian Global Significance.
@@
Table 6: Ablation with varying SH compactness: With a Gaussian pruning ratio of 66% and SH reduced to 2 degrees, LightGaussian nearly maintains rendering quality (SSIM drops slightly from 0.927 to 0.926). Reducing SH further to 1 degree lowers SSIM to 0.923.
Figure 6: Rendering performance comparison across different pruning levels and VQ ratios. Note that in subfigure (b), we use the VQ ratio to indicate the proportion of SH parameters quantized to an 8192-codebook, while the (overall) quantization ratio reflects combined compression from VQ and FP32-to-FP16 bitwise quantization.
@@
5 Conclusion, Limitations, and Broad Impact
We present LightGaussian, a novel framework that transforms heavy point-based representations into a compact format for efficient novel view synthesis. Designed for practical use, LightGaussian leverages 3D Gaussians to model large-scale scenes, effectively identifying and pruning the least significant Gaussians generated through densification. It achieves over 15$\times$ data reduction, boosts FPS to over 200, and minimally impacts rendering quality. Exploring zero-shot compression across various 3D-GS-based frameworks remains a promising direction for future research.
Broadly, LightGaussian’s compact representation has the potential to democratize high-quality 3D content for applications in VR, AR, and autonomous driving by reducing resource demands, enabling more accessible and scalable deployment across industries. While we do not foresee any significant risk from this specific technique, 3D reconstruction may infringe on personal privacy when applied in public spaces or with drone footage, contradicting our ethical intentions.
Appendix